1. Show by using an (ε − N ) argument that lim n->inf 3n^2 −2n+1/2n^2 − 4 =3/2
  2. Use an (ε − δ) argument to show that f : R → R be the function defined by

f(x)= {x^2−5x−5 if x≥−1 x^2+x+1 if x<−1}

is continuous at x = −1


1
Expert's answer
2021-05-12T05:07:29-0400

1.)

"x_n={3n^2-2n+1 \over 2n^2-4 }"

"|x_n-\frac{3}{2} |=|\frac{3n^2-2n+1}{2n^2-4}-\frac{3}{2}|"


"=|\frac{6n^2-4n+2-(6n^2-12)}{2(2n^2-4)}|"


"|x_n-\frac{3}{2} |=\frac{|4n-7|}{|4n^2-8|}=\frac{2n-7}{2n^2-4}" for "n\ge 4"


now

"|x_n-\frac{3}{2} |=\frac{2n-7}{2n^2-4} <\frac{2n}{2n^2-4}=\frac{n}{n^2-2}"


but "n^2-2>\frac{n^2}{2} \ for \ n \ge4"


thus

"|x_n-\frac{3}{2} |<\frac{2n}{n^2}= \frac{2}{n}<\epsilon \ \ \ \ \ if \ n> \frac{2}{\epsilon}"


thus

"|x_n-\frac{3}{2} |=\epsilon \ \ \ \ \ \ \ \ \ for \ all \ n \ge [\frac{2}{\epsilon},4]"



2.)

"f(x) = \begin{cases}
 x^2-5x-5 \ \ \ \ \ & x \ge-1 \\
 x^2+x+1 & x\le-1
\end{cases}"


"f(-1)=1"


let choose "\delta" such that "0<\delta <1"

thus "|x-1|<\delta \ \ \ \ \ \ \implies |x+1|<1"

thus "|x-6|=|x+1-7|<|x+1|+7<8"

and "|x|=|x+1-1|<|x+1|+1<2<8"


now for "0<\delta <1" , we have "|x-6|<8 \ , |x|<\delta"

thus "|f(x)-1|<|x+1|.\delta < \epsilon \ \ if \ |x+1|<\frac{\epsilon}{\delta}"


hence choose "\delta" = min {"\frac{\epsilon}{\delta}," 1 } , we get


"|f(x)-1|<\epsilon" for all "|x+1|< \delta"

and "\delta = min" {"\frac{\epsilon}{\delta}," 1}